A company had the following purchases and sales during its first year of operations:


PurchasesSales
January:10 units at $1206 units
February:20 units at $1255 units
May:15 units at $1309 units
September:12 units at $1358 units
November10 units at $14013 units

On December 31, there were 26 units remaining in ending inventory. Using the PERPETUAL LIFO inventory valuation method, what is the cost of the ending inventory? (Assume all sales were made on the last day of the month.)

A) $3,405

B)$3,270

C)$3,364

D)$5,400

On December 31, there were 26 units remaining in ending inventory. Using the PERIODIC LIFO inventory valuation method, what is the cost of the ending inventory? (Assume all sales were made on the last day of the month.)

A)$3,405

B)$3,270

C)$3,200

D)$3,364

E)$5,400

On December 31, there were 26 units remaining in ending inventory. Using the PERPETUAL FIFO inventory valuation method, what is the value of cost of goods sold? (Assume all sales were made on the last day of the month.)

A)$8,670

B)$3,540

C)$5,400

D)$5,130

E)$3,270

On December 31, there were 26 units remaining in ending inventory. Using the PERPETUAL LIFO inventory valuation method, what is the value of cost of goods sold? (Assume all sales were made on the last day of the month.)

A)$8,670

B)$5,400

C)$5,470

D)$5,130

E)$5,305

On December 31, there were 26 units remaining in ending inventory. Using the PERIODIC LIFO inventory valuation method, what is the value of cost of goods sold? (Assume all sales were made on the last day of the month.)

A)$8,670

B)$5,400

C)$5,470

D)$3,200

E)$5,130
Answer #1

Requirement 1 - Option (B) is Correct LIFO Method Perpetual inventory Record Date Purchases Cost of goods sold Quantity unitRequirement 4 - Option (B) is Correct LIFO Method Perpetual inventory Record Date Purchases Cost of goods sold Quantity unitPeriodic Method Requirement 2 - Option [C] is correct LIFO Method: Activity Jan purchase Feb purchase May purchase purchase pPeriodic Method Requirement 5 - Option [C] is correct LIFO Method: Activity Jan purchase Feb purchase May purchase purchase pRequirement 3 - Option (D) is Correct FIFO Method Perpetual inventory Record Date Purchases Cost of goods sold Quantity unit
Answer #3

Ans: (D) $3540

Explanation:

Nov. 10 units * $140 $1400
Sep. 12 units * $135 $ 1620
May (26- 10-12)= 4 units * $130 $520
Total 26 Unit $3540
Answer #2
Cost of the ending inventory = (10*120)+(16*125)= $3200